# A Generalized Information Bottleneck Theory of Deep Learning

### v2

@@ 1 Introduction
Deep learning has achieved remarkable practical success, yet our theoretical understanding of how neural networks learn effective representations remains incomplete (Shwartz-Ziv & Tishby, 2017). Information theory offers a principled framework for analyzing deep learning, as information-theoretic quantities are invariant to invertible transformations and provide interpretable units of measurement (Cover & Thomas, 1991). The Information Bottleneck (IB) principle, introduced by Tishby et al. (1999), has emerged as a particularly influential framework for understanding neural network learning dynamics, providing
@@
insights into diverse phenomena including adversarial robustness (Ma et al., 2021), the effects of dropout regularization (Achille & Soatto, 2018), and generalization bounds (Kawaguchi et al., 2023). Through this interpretation, the activations of our network can be viewed as a hidden state representation $\mathcal{T}$, which converges to a set of statistics captured by two competing terms. The first term, referred to as the prediction term, quantifies the mutual information (MI) between the hidden representation $\mathcal{T}$ and the target $Y$, denoted as $I(Y;\mathcal{T})$ [^1] For the mathematical notation used throughout this paper, see Appendix A.. It is straightforward to see that achieving training (and consequently test) accuracy above random guessing requires a network whose learned representation is well aligned with that of the target data. However, it is well established that optimizing solely for prediction accuracy can lead to overfitting. Consequently, the IB framework introduces a second term: the complexity term $I(\mathcal{X};\mathcal{T})$, which quantifies the mutual information between the input data $\mathcal{X}$ and the hidden representation. Optimizing this less intuitive term can be interpreted as an effort to minimize redundant and irrelevant information from the input that is encoded in the latent space.
@@
Shwartz-Ziv & Tishby (2017) suggested that the success of deep learning can be attributed to the ability of NNs to perform the aforementioned optimization problem in two distinct phases. First, a fitting phase, where both of the two introduced terms increase, and second a compression phase, where the complexity term decreases in size (refer to Figure 1 for a visual illustration of these processes). It was argued that this second compression phase was unique to deep models and helped explain their generalizability. In some cases, the flow of information through the latent space has been shown to align precisely with the IB’s theoretical predictions (Shwartz-Ziv & Tishby, 2017).
While the IB framework initially seemed to provide a complete explanation of how neural networks balance compression and predictive accuracy, Saxe et al. (2019) presented counterexamples that challenge this view. In particular, the authors showed that compression phases depend critically on the choice of activation function: networks with tanh activations exhibited compression across all layers, whereas ReLU-based networks did not. Despite the absence of a compression phase, the ReLU networks still generalized well. According to Goldfeld et al. (2019), this occurs because the complexity term in deterministic networks is theoretically constant or infinite, rendering compression impossible. Consequently, the compression observed in tanh networks was not a genuine information-theoretic effect, but rather the result of injected randomness (Saxe et al., 2019; Shwartz-Ziv & Tishby, 2017; Geiger, 2022).
In this paper, we address these issues by introducing a generalized formulation of the IB framework that is grounded in synergy. Synergy, a concept from multivariate information theory, captures the extra predictive power that arises when inputs are considered together rather than in isolation (Williams & Beer, 2010). To motivate this perspective, we begin by asking: why synergy? We then present both theoretical arguments and empirical results showing that synergistic functions lead to improved generalization.
@@
Figure 2: Information plane dynamics across multiple activation functions, extending Shwartz-Ziv & Tishby (2017) and Saxe et al. (2019) beyond tanh and ReLU to include softplus, swish, and leaky ReLU. Standard IB (blue) shows compression only for tanh; GIB (pink) shows compression for all activation functions. Each column represents one seed.
@@ 2 Related Work
MI estimation in the IB framework remains an active area of research and debate. Numerous estimators exist (e.g., $k$-nearest-neighbor (kNN) methods (Kozachenko & Leonenko, 1987; Kraskov et al., 2004), and kernel-density approaches (Kandasamy et al., 2015; Han et al., 2017)), and trainable neural estimators (Belghazi et al., 2018)), yet many information-theoretic studies of deep networks (Shwartz-Ziv & Tishby, 2017; Saxe et al., 2019) discretize neuron outputs (“binning”) to approximate MI. Binning is simple and fast, but even moderate coarse-graining can introduce substantial estimation error (Goldfeld et al., 2019). Despite these limitations, we use binning because MI estimates are needed throughout training (e.g., at each epoch); running kNN, KDE, or variational estimators at this frequency would be prohibitively slow and numerically unstable in high-dimensional settings. Our goal is to track relative trends in MI rather than obtain exact differential values, and binning provides a tractable, reproducible proxy that makes per-training-step MI monitoring feasible.
@@
There exist many extensions of the IB framework that move beyond the original two-variable setting. For instance, the Variational IB introduces stochastic neural parameterizations to scale IB to deep networks (Alemi et al., 2017), while Information Dropout applies a neuron-wise IB-like penalty to improve generalization (Achille & Soatto, 2018). More recent work has drawn connections between IB and the framework of Partial Information Decomposition (PID), with Kolchinsky et al. (2019); Kolchinsky (2024) showing that redundancy can be isolated via an IB-style optimization. However, the work most closely related to ours is the multivariate IB of Friedman et al. (2001), which explicitly captures structure within the latent space by introducing multiple bottleneck variables and using graphical models to specify both compression and preservation relations among them. This enables the IB to consider how inputs interact to describe latent representations. While Friedman et al. (2001) multivariate IB captures statistical dependencies between multiple bottleneck variables, it does not explicitly quantify or optimize for synergistic information processing. Our approach fundamentally differs by directly measuring the information available only through collective feature processing, which we posit is key to understanding generalization in deep networks.
@@
Synergy characterizes the additional information obtained by evaluating variables collectively rather than individually, quantifying how features interact to reduce uncertainty about a target. The characteristics of this relationship can be illustrated by means of the XOR function. Consider two binary string variables, $X_{1}$ and $X_{2}$, with $Z$ being their XOR output. In this scenario, $X_{1}$ and $Z$, as well as $X_{2}$ and $Z$, are uncorrelated ($I(X_{1};Z)=I(X_{2};Z)=0$), but together, $X_{1}$ and $X_{2}$ fully describe $Z$ ($I(X_{1},X_{2};Z)=H(Z)$) (Guyon & Elisseeff, 2003; Williams & Beer, 2010). While the concept is intuitive, its formalization has proven challenging, leading to multiple proposed measures. Early work by Williams & Beer (2010) introduced PID, which decomposes MI into unique, redundant, and synergistic components. However, the number of terms in this decomposition equals the n1n-1’th Dedekind number, where $n$ is the number of features. This number is impractically large: a system with nine variables would require approximately $5\times 10^{22}$ terms, while for ten variables, the Dedekind number remains unknown. Moreover, estimating these terms is subject to convergence issues and size limitations (Makkeh et al., 2018; 2019; Pakman et al., 2021). While Varley & Hoel (2022) reduced the number of investigable quantities by averaging contributions of layers in the PID lattice, with different layers representing different levels of redundancy or synergy, calculations remained too complex for applications involving more than a few features. Alternative measures such as O-information (Rosas et al., 20019), correlational importance (Nirenberg & Latham, 2003), and synergistic MI (Griffith & Koch, 2014) can estimate the synergy or redundancy of large sets of variables, but fail to reveal whether a specific feature interacts synergistically or redundantly. We resolve these issues by averaging the interaction information of each feature with those remaining:
@@ 3.1.1 Theoretical Evidence
This quantity can be re-written as the complexity term of the IB, reducing its value has repeatedly been shown to be related to compression and generalization (Tishby et al., 1999; Shwartz-Ziv & Tishby, 2017). High values of this term ensure a latent representation that has memorized irrelevant and redundant information in the input. Furthermore, recent work has formally related this quantity to generalization bounds (Kawaguchi et al., 2023).
@@ 3.2 Formulating the GIB Principle
To facilitate this, we define a new distribution that, instead of describing $Z$ our output, describes the co-occurrences of $Z$ with $Y$. The exact definition of $Q(Y,Z)$ is based on PMI-based reweighting, i.e., weighting samples by the likelihood ratio between the joint distribution and the product of marginals: $Q(Z,Y)=\frac{P(Z,Y)}{P(Z)P(Y)}$. This reweighting scheme emphasizes patterns where $Z$ and $Y$ co-occur more frequently than would be expected under independence, effectively highlighting the meaningful dependencies between our learned representations and the target outputs. PMI has proven effective in capturing meaningful associations in numerous ML contexts: it underlies word2vec’s implicit matrix factorization (Levy & Goldberg, 2014), drives contrastive learning objectives (van den Oord et al., 2018), and measures feature relevance in interpretable ML (Bouma, 2009). Combining this definition with how we earlier defined synergy, we get the following formulation of the GIB:
$\mathcal{L}_{\text{GIB}}$ $=\max_{p(Z|X)}\left[\text{Syn}(\mathcal{X}\to Q(Z,Y))\right]$ (5)
$=\max_{p(Z|X)}\left[\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}I(\mathcal{X};Q(Z,Y))}}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{prediction term}}}-\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Q(Z,Y))+I(X^{i});Q(Z,Y))\right)}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{complexity term}}}\right]$ (6)
The prediction term (blue) $I(\mathcal{X};Q(Z,Y))$ measures the mutual information between all input features and the PMI-reweighted distribution, capturing how well the complete feature set predicts patterns where model outputs and true labels co-occur beyond chance. The complexity term (red) $\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Q(Z,Y))+I(X^{i};Q(Z,Y))\right)$ inversely quantifies the average information obtainable from individual features or their complements. By maximizing their difference, GIB measures information dynamics that emerge only from collective feature interactions, which our analysis also indicates leads to improved generalization. On the other hand, measuring synergy can be computationally demanding, as we discuss in Appendix G.
@@ 4 Relating the GIB to the IB
If we assume perfect training accuracy and therefore $Q(Z,Y)=Z=Y$, then the original IB objective is upper bounded by our GIB:
$I(\mathcal{T};Y)-\beta I(\mathcal{X};\mathcal{T})\leq I(\mathcal{X};Q(Z,Y))-\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Q(Z,Y))+I(X^{i};Q(Z,Y))\right)$ (7)
@@
This new formulation overcomes two main limitations of the original IB. First, the partition across subsets of features combined with the PMI definition of $Q(Z,Y)$ protects the compression term from becoming infinite. In Appendix D we prove that the GIB is only infinite under interpretable circumstances. Second, and more fundamentally, our formulation explicitly considers over-reliance on individual features, as explained in Section 3.1. IB optimizes the total information flow between inputs and outputs through the latent representation $\mathcal{T}$ without considering how features interact. In contrast, GIB explicitly models how inputs combine to form the latent representation, distinguishing between different types of feature interactions. This is evident in the complexity terms: IB’s $\beta^{-1}I(\mathcal{X};\mathcal{T})$ aggregates all information equally, while GIB’s synergistic decomposition $\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Q(Z,Y))+I(X^{i};Q(Z,Y))\right)$ penalizes the information contained in individual features. Consequently, IB compresses indiscriminately, whereas GIB selectively preserves long-range feature combinations; the synergistic patterns we have shown lead to better generalization.
@@ 5.2 Information Dynamics of ResNets
We analyze information dynamics in residual networks (ResNets) of varying depths (20, 56, 80, 110 layers) trained on CIFAR-10. For the standard IB, we compute MI using the 10-dimensional output layer directly. For GIB, due to the need to compute feature-wise decompositions on the high-dimensional input space (3072 dimensions), we first apply Kernel PCA to reduce the pixel space to 50 principal components before computing MI, as explained in more detail in Appendix E.4 (Turk & Pentland, 1991). For more details as to why we chose 50 components, see Appendix H.
Figure 5 displays information plane trajectories for ResNets trained on CIFAR-10. The GIB formulation shows consistent compression behavior across all network depths and random seeds, though the dynamics vary with architecture size. In smaller networks (ResNet-20), trajectories show a general trend of increasing prediction term while the complexity term decreases throughout training. Larger networks (ResNet-56 and above) begin to show more pronounced phase structure.
The standard IB presents markedly different dynamics. Rather than showing clear phases, IB trajectories remain largely clustered with minimal directional movement across epochs. The absence of compression phases in standard IB holds across all tested architectures, confirming previous observations that ReLU networks fail to exhibit expected IB behavior (Saxe et al., 2018).

### v3

@@ 1 Introduction
Deep learning has achieved remarkable practical success, yet our theoretical understanding of how neural networks learn effective representations remains incomplete (Shwartz-Ziv and Tishby, 2017). Information theory offers a principled framework for analyzing deep learning, as information-theoretic quantities are invariant to invertible transformations and provide interpretable units of measurement (Cover and Thomas, 1991).
@@
The Information Bottleneck (IB) principle, introduced by Tishby et al. (1999), has emerged as a particularly influential framework for understanding neural network learning dynamics, providing insights into diverse phenomena including adversarial robustness (Ma et al., 2021), the effects of dropout regularization (Achille and Soatto, 2018), and generalization bounds (Kawaguchi et al., 2023). Through this interpretation, the activations of our network can be viewed as a hidden state representation $\mathcal{T}$, which converges to a set of statistics captured by two competing terms. The first term, referred to as the prediction term, quantifies the mutual information (MI) between the hidden representation $\mathcal{T}$ and the target $Y$, denoted as $I(Y;\mathcal{T})$ [^1] For the mathematical notation used throughout this paper, see Appendix A.. It is straightforward to see that achieving training (and consequently test) accuracy above random guessing requires a network whose learned representation is well aligned with that of the target data. However, it is well established that optimizing solely for prediction accuracy can lead to overfitting. Consequently, the IB framework introduces a second term: the complexity term $I(\mathcal{X};\mathcal{T})$, which quantifies the mutual information between the input data $\mathcal{X}$ and the hidden representation. Optimizing this less intuitive term can be interpreted as an effort to minimize redundant and irrelevant information from the input that is encoded in the latent space.
@@
Shwartz-Ziv and Tishby (2017) suggested that the success of deep learning can be attributed to the ability of NNs to perform the aforementioned optimization problem in two distinct phases. First, a fitting phase, where both of the two introduced terms increase, and second a compression phase, where the complexity term decreases in size (refer to Figure 1 for a visual illustration of these processes). It was argued that this second compression phase was unique to deep models and helped explain their generalizability. In some cases, the flow of information through the latent space has been shown to align precisely with the IB’s theoretical predictions (Shwartz-Ziv and Tishby, 2017).
While the IB framework initially seemed to provide a complete explanation of how neural networks balance compression and predictive accuracy, Saxe et al. (2019) presented counterexamples that challenge this view. In particular, the authors showed that compression phases depend critically on the choice of activation function: networks with tanh activations exhibited compression across all layers, whereas ReLU-based networks did not. Despite the absence of a compression phase, the ReLU networks still generalized well. According to Goldfeld et al. (2019), this occurs because the complexity term in deterministic networks is theoretically constant or infinite, rendering compression impossible. Consequently, the compression observed in tanh networks was not a genuine information-theoretic effect, but rather the result of injected randomness (Saxe et al., 2019; Shwartz-Ziv and Tishby, 2017; Geiger, 2022).
In this paper, we address these issues by introducing a generalized formulation of the IB framework that is grounded in synergy. Synergy, a concept from multivariate information theory, captures the extra predictive power that arises when inputs are considered together rather than in isolation (Williams and Beer, 2010). To motivate this perspective, we begin by asking: why synergy? We then present both theoretical arguments and empirical results showing that synergistic functions lead to improved generalization.
@@
Figure 2: Information plane dynamics across multiple activation functions, extending Shwartz-Ziv and Tishby (2017) and Saxe et al. (2019) beyond tanh and ReLU to include softplus, swish, and leaky ReLU. Standard IB (blue) shows compression only for tanh; GIB (pink) shows compression for all activation functions. Each column represents one seed.
@@ 2 Related Work
MI estimation in the IB framework remains an active area of research and debate. Numerous estimators exist (e.g., $k$-nearest-neighbor (kNN) methods (Kozachenko and Leonenko, 1987; Kraskov et al., 2004), and kernel-density approaches (Kandasamy et al., 2015)), and trainable neural estimators (Belghazi et al., 2018)), yet many information-theoretic studies of deep networks (Shwartz-Ziv and Tishby, 2017; Saxe et al., 2019) discretize neuron outputs (“binning”) to approximate MI. Binning is simple and fast, but even moderate coarse-graining can introduce substantial estimation error (Goldfeld et al., 2019). Despite these limitations, we use binning because MI estimates are needed throughout training (e.g., at each epoch); running kNN, KDE, or variational estimators at this frequency would be prohibitively slow and numerically unstable in high-dimensional settings. Our goal is to track relative trends in MI rather than obtain exact differential values, and binning provides a tractable, reproducible proxy that makes per-training-step MI monitoring feasible.
@@
There exist many extensions of the IB framework. For instance, the variational information bottleneck (VIB) introduces stochastic neural parameterizations to scale IB to deep networks (Alemi et al., 2017), while Information dropout applies a neuron-wise IB-like penalty via multiplicative noise to improve generalization (Achille and Soatto, 2018). The HSIC bottleneck (Wang et al., 2021) replaces mutual information with the Hilbert–Schmidt independence criterion to regularize intermediate representations for adversarial robustness. More recent work has extended the IB to multivariate and deep settings. Matrix-based Rényi’s $\alpha$-order entropy functionals (Yu et al., 2020; 2021a) provide estimators of multivariate entropies and dependence. These have been used to develop deterministic IB objectives in deep networks (Yu et al., 2021b), and to design gated IB objectives for sequential environments (Alesiani et al., 2023). The multivariate IB of Friedman et al. (2001) captures statistical dependencies between multiple bottleneck variables through graphical models. Most relevant to our work, recent advances have also connected IB and Partial Information Decomposition (PID), with Kolchinsky (2024) building on earlier IB work such as Kolchinsky et al. (2019), showing that PID redundancy can be isolated via an IB-style optimization. Although these methods extend IB to multivariate settings and even show that certain PID quantities, such as redundancy, can be characterized via bottlenecks, none explicitly incorporate a synergy-specific information term into the IB formulation. Our approach achieves this, and we will demonstrate that it not only results in more consistent compression behavior but also offers theoretical advantages.
@@
Synergy characterizes the additional information obtained by evaluating variables collectively rather than individually, quantifying how features interact to reduce uncertainty about a target. The characteristics of this relationship can be illustrated by means of the XOR function. Consider two binary string variables, $X_{1}$ and $X_{2}$, with $Z$ being their XOR output. In this scenario, $X_{1}$ and $Z$, as well as $X_{2}$ and $Z$, are uncorrelated ($I(X_{1};Z)=I(X_{2};Z)=0$), but together, $X_{1}$ and $X_{2}$ fully describe $Z$ ($I(X_{1},X_{2};Z)=H(Z)$) (Guyon and Elisseeff, 2003; Williams and Beer, 2010). While the concept is intuitive, its formalization has proven challenging, leading to multiple proposed measures. Early work by Williams and Beer (2010) introduced PID, which decomposes MI into unique, redundant, and synergistic components. However, the number of terms in this decomposition equals the n1n-1’th Dedekind number, where $n$ is the number of features. This number is impractically large: a system with nine variables would require approximately $5\times 10^{22}$ terms, while for ten variables, the Dedekind number remains unknown. Moreover, estimating these terms is subject to convergence issues and size limitations (Makkeh et al., 2018; 2019; Pakman et al., 2021). While Varley and Hoel (2022) reduced the number of investigable quantities by averaging contributions of layers in the PID lattice, with different layers representing different levels of redundancy or synergy, calculations remained too complex for applications involving more than a few features. Alternative measures such as O-information (Rosas et al., 2019), correlational importance (Nirenberg and Latham, 2003), and synergistic MI (Griffith and Koch, 2014) can estimate the synergy or redundancy of large sets of variables, but fail to reveal whether a specific feature interacts synergistically or redundantly. We resolve these issues by averaging the interaction information of each feature with those remaining:
@@ 3.1.1 Theoretical Evidence
This quantity can be re-written as the complexity term of the IB, reducing its value has repeatedly been shown to be related to compression and generalization (Tishby et al., 1999; Shwartz-Ziv and Tishby, 2017). High values of this term ensure a latent representation that has memorized irrelevant and redundant information in the input. Furthermore, recent work has formally related this quantity to generalization bounds (Kawaguchi et al., 2023).
@@ 3.2 Formulating the GIB Principle
To facilitate this, we take the following two steps. First, our prediction term will solely measure the MI between our predictions and targets. Second, our complexity term will be a function of a new distribution that describes the co-occurrences of $Z$ with $Y$. The exact definition of $Q$ is based on PMI-based reweighting, i.e., weighting samples by the likelihood ratio between the joint distribution and the product of marginals: $Q(Z,Y)=\frac{P(Z,Y)}{P(Z)P(Y)}$. This reweighting scheme emphasizes patterns where $Z$ and $Y$ co-occur more frequently than would be expected under independence, effectively highlighting the meaningful dependencies between our learned representations and the target outputs. PMI has proven effective in capturing meaningful associations in numerous ML contexts: it underlies word2vec’s implicit matrix factorization (Levy and Goldberg, 2014), drives contrastive learning objectives (van den Oord et al., 2018), and measures feature relevance in interpretable ML (Bouma, 2009). By sampling from the distribution $Q(Z,Y)=\frac{P(Z,Y)}{P(Z)P(Y)}$ we obtain the random variable $Q$. Combining these steps with how we earlier defined synergy, we get the following formulation of the GIB:
$\mathcal{L}_{\text{GIB}}$ $=\max_{p(Z|X)}\left[\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}I(Z;Y)}}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{prediction term}}}-\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Q)+I(X^{i};Q)\right)}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{complexity term}}}\right]$ (5)
The prediction term (blue) $I(Z;Y)$ measures the mutual information between the model outputs $Z$ and the labels $Y$, capturing how well the predictions align with the true targets. The complexity term (red) $\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Q)+I(X^{i};Q)\right)$ inversely quantifies the average information obtainable from individual features or their complements about the PMI-reweighted distribution $Q(Z,Y)$, which emphasizes correct predictions. By maximizing their difference, GIB measures information dynamics that emerge only from collective feature interactions, which our analysis also indicates leads to improved generalization. On the other hand, measuring synergy can be computationally demanding, as we discuss in Appendix G.
@@ 4 Relating the GIB to the IB
If we assume perfect training accuracy and therefore $Q=Z=Y$, then the original IB objective is upper bounded by our GIB:
$I(\mathcal{T};Y)-\beta I(\mathcal{X};\mathcal{T})\leq I(Z;Y)-\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Q))+I(X^{i};Q))\right)$ (6)
@@
This new formulation overcomes two main limitations of the original IB. First, the partition across subsets of features combined with the PMI definition of $Q(Z,Y)$ protects the compression term from becoming infinite. In Appendix D we prove that the GIB is only infinite under interpretable circumstances. Second, and more fundamentally, our formulation explicitly considers over-reliance on individual features, as explained in Section 3.1. IB optimizes the total information flow between inputs and outputs through the latent representation $\mathcal{T}$ without considering how features interact. In contrast, GIB explicitly models how inputs combine to form the latent representation, distinguishing between different types of feature interactions. This is evident in the complexity terms: IB’s $\beta^{-1}I(\mathcal{X};\mathcal{T})$ aggregates all information equally, while GIB’s synergistic decomposition $\frac{1}{2\beta N}\sum_{i=1}^{N}\left(I(\mathcal{X}^{-i};Q)+I(X^{i};Q)\right)$ penalizes the information contained in individual features. Consequently, IB compresses indiscriminately, whereas GIB selectively preserves long-range feature combinations; the synergistic patterns we have shown lead to better generalization.
@@ 5.2 Information Dynamics of ResNets
We analyze information dynamics in residual networks (ResNets) of varying depths (20, 56, 80, 110 layers) trained on CIFAR-10. For the standard IB, we compute MI using the 10-dimensional output layer directly. For GIB, due to the need to compute feature-wise decompositions on the high-dimensional input space (3072 dimensions), we first apply Kernel PCA to reduce the pixel space to 50 principal components before computing MI, as explained in more detail in Appendix E.4 (Turk and Pentland, 1991). For more details as to why we chose 50 components, see Appendix H.
Figure 5 displays information plane trajectories for ResNets trained on CIFAR-10. The GIB formulation shows consistent compression and fitting for all network depths and random seeds, though the dynamics vary with architecture size. In smaller networks (ResNet-20), trajectories show a general trend of increasing prediction term while the complexity term decreases throughout training. Larger networks (ResNet-56 and above) begin to show phase structure.
The standard IB presents markedly different dynamics. Rather than showing clear phases, IB trajectories remain largely clustered with minimal compression across epochs. The absence of compression phases in standard IB holds across all tested architectures, confirming previous observations that ReLU networks fail to exhibit expected IB behavior (Saxe et al., 2018).